OpenStudy (anonymous):

consider three points, (3,7),(-6,1) and (9,p)on the same line. find value of p and explain the steps.

OpenStudy (anonymous):

11

OpenStudy (turingtest):

Find the equation of this line. Fist find the slope: \[m={y_2-y_1 \over x_2-x_1}={1-7 \over -6-3}=2/3\]Then use point slope form:\[y-y_1=m(x-x_1)\]\[y-1=(2/3)(x+6)=2x/3+4\]\[y=2x/3+5\]mow we can find y for any given x. the given x in the question is 9 so \[y=2(9)/3+5=6+5=11=p\]

OpenStudy (anonymous):

hi turing how to draw didvide y line for several components

OpenStudy (turingtest):

what do you mean? the horizontal divide lines?

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

If they are on the same line that means they have the same slope. So you'll use the first two points to find the slope of the line using the slope formula which says the difference of the y's divided by the difference of the x's: (1-7)/(-6-3)=-6/-9=2/3. Now that you have the slope (2/3), you can find the equation of the line using this formula: y=mx+b--m being the slope 2/3 and then you use one of the points to find b. I'm going to use (3,7) 7=(2/3)3+b 7=2+b 5=b So, the equation of the line is y=(2/3)x + 5 Now that you can plug in 9 into x to find y in your case p y=(2/3)9+5 y=(2/3)9+5 y=11

OpenStudy (turingtest):

you write over instead of a division symbol and enclose it in brackets: {x over y} \[{x \over y}\]make sure to put spaces between letters and the word "over"

OpenStudy (turingtest):

the brackets control where the line starts and stops

OpenStudy (anonymous):

{x - y over 2}

OpenStudy (turingtest):

gotta put it in the equation feature. Highlight it and select "equation"

OpenStudy (turingtest):

\[{x - y \over 2}\]

OpenStudy (anonymous):

thanks got that \[{{y - 5 \over 5} +2 \over 7}\]

OpenStudy (turingtest):

nice, never tried that but thought it was possible :)